OpenStudy (anonymous):

A rocket is fired straight up from a 60 ft. platform with an initial velocity of 96 ft/sec. The height of the rocket, h(t), is found using the function h(t) = - 16t2 + 96t + 60 where t is the time in seconds. Find the maximum height. 190 ft 7 ft 204 ft 202 ft

OpenStudy (danjs):

the physics really is just extra info they just want to know the maximum value for the function h(t)

OpenStudy (boldjon):

h(t) = -16t^2 + 96t + 60 v(t) = -32t + 96 at max height , v = 0 -32t + 96=0 32t = 96 t = 96/32 = 3 h(3) = -16(9) + 96(3) + 60 = 204 or h(t) = -16(t^2 - 6t + 9-9) + 60 = -16( (t-3)^2 - 9) + 60 = -16(t-3)^2 + 144 + 60 = -16(t-3)^2 + 204 this has a vertex of (3,204) thus a max of 204 when t - 3

OpenStudy (danjs):

yeah the derivative works to figure when the velocity is zero at the maximum point

OpenStudy (raffle_snaffle):

or show him the derivation for the kinematic equations

OpenStudy (boldjon):

so i was right? 204?

OpenStudy (danjs):

you said for algebra though, the last time you just remembered the axis of symmetry thing t = -B/(2A) = -96/(2*-16) = 3 the maximum value for that flipped parabola is on that symmetry line, and is the vertex point h(3)